OpenStudy (kanwal32):

how find gamma of -15/2

OpenStudy (kanwal32):

@tHe_FiZiCx99

OpenStudy (kainui):

There are a handful of ways you can do this, depending on what you wanna do. You can remember this value (or derive it from the integral definition with a cute trick using the Gaussian) \[\Gamma(\tfrac{1}{2}) = \sqrt{\pi}\] and then remember this (or derive it from the integral definition using integration by parts) \[\Gamma(n+1)=n\Gamma(n)\] That's really just the Gamma version of \(n!=n*(n-1)!\) but keep in mind that the Gamma version works for non integers!

OpenStudy (kainui):

So now you can use this recursive definition to get \(\Gamma( \tfrac{-15}{2})\) Here's the first step getting you closer to \(\Gamma(\tfrac{1}{2})\) so that you can exactly evaluate it: \[ \Gamma( \tfrac{-13}{2}) = \tfrac{-15}{2}\Gamma( \tfrac{-15}{2}) \] See if you can do this quickly rather than multiplying each and every piece.